OpenStudy (anonymous):

If i get the correct answer i will medal and fan. I don't understand how to do this at all. Alex stocks up for winter. He buys 32 cans of vegetables. He pays 80 cents per can for tomatoes and 40 cents per can for corn, for a total cost of $18. How many cans of tomatoes does he buy?

OpenStudy (anonymous):

@princeharryyy @natsent06

OpenStudy (anonymous):

I think you need to set up a system of equations

OpenStudy (anonymous):

.8t + .4c = 18 t + c = 32

OpenStudy (anonymous):

let cans of corn be x let cans of tomatoes be y x+y = 32 ( as alex have 32 cans of vegtables) 40x +80y = 1800 cents Find x and y

OpenStudy (anonymous):

solve by substitution or elimination

OpenStudy (princeharryyy):

13